OpenStudy (anonymous):

how can I write 1*3*5*7...??

OpenStudy (across):

That could be interpreted two ways: 1. as the product of consecutive primes, or 2. as the product of consecutive odd integers.

OpenStudy (across):

;P

OpenStudy (turingtest):

we could be fancy and write it with the \(\Pi\) notation :)

OpenStudy (across):

If it's the latter, then\[\prod_{i=0}^{\infty}2i+1\]

OpenStudy (turingtest):

like that^

OpenStudy (anonymous):

is as a product of odd numbers

OpenStudy (across):

Well, if it was the former, and you had an answer, then I would suggest that we get together. ;P

OpenStudy (anonymous):

is just I have to write it in a general way is for a differential equation

OpenStudy (turingtest):

but across, 1 is not prime and 2 is not on the list so it must be option 2, nno?

OpenStudy (across):

I am just messing around. ^^

OpenStudy (anonymous):

A minor problem is that 1*3*5*7*... is meaningless.

OpenStudy (anonymous):

Are you wondering how to express the first n terms? That would be (2*1-1)(2*2-1)(2*3-1)...(2*n-1)

OpenStudy (anonymous):

yes just like that the thing is that I'm wondering is possible that I could write that like a factorial?

OpenStudy (anonymous):

Not one factorial. But compare just the format through 7, with 7! and you have 1*2*3*4*5*6*7 as the factorial and 1*3*5*7 as the number you want. How do you get from the first to the second?

OpenStudy (anonymous):

What happens when you divide the factorial by 2*4*6?

OpenStudy (anonymous):

my equation is (x-1)y''-xy'+y=0 in my solution I have to state 1*3*5*7...in the series

OpenStudy (anonymous):

Notice that 1*3*5*7 is 1*2*3*4*5*6*7 divided by 2*4*6. Both of these can be expressed using factorials.

OpenStudy (anonymous):

I don't like just giving answers, but gotta run. The top is (2n+1)! for and the bottom is 2*n! for n=3.

OpenStudy (anonymous):

Oh! you're si right!!! but how can you state that just in factorials? sorry is just I can't see how U_U

OpenStudy (anonymous):

The nth term in your sequence is (2n+1)! / (2*n!)